OpenStudy (anonymous):

Please help!!! Let (-5,2) be a point on the terminal side of theta find the exact values of cos theta, csc theta, and tan theta

OpenStudy (er.mohd.amir):

not possible -1<=cos x<=1

OpenStudy (nayef):

Sorry: cos(θ)=-\[\frac{ 5 }{ \sqrt{29} }\], , \[\csc(\theta)=\frac{ \sqrt{29} }{ 2} , \tan(\theta)=-2/5\]

OpenStudy (nayef):

Be careful! there a minus in the cos value!

OpenStudy (er.mohd.amir):

only cosec +ve and cos and tan -ve answers are correct by "Nayel"

OpenStudy (er.mohd.amir):

sorry Nayef

OpenStudy (nayef):

Welcome, It is a great chance to learn from my mistakes.